# SCHEME: Scalable Channer Mixer for Vision Transformers

### v1

@@ 4.2 Ablation Studies
In this section, we discuss ablations of the contribution of the BD-MLP and CCA branches, alternative designs for the channel mixer, and the regularizing effect of CCA.
@@
Contribution of BD-MLP and CCA branch: Table 6 shows an ablation of the contribution of the BD MLP and CCA branches to transformer performance. Starting from the Metaformer-PPAA-11-e2-S12, with expansion ratio $E=2$ and dense MLP, we replace the channel mixer with that of the SCHEME to obtain SCHEMEformer-PPAA-44-e8-S12, which has $g_{1}=g_{2}=4$ and $E=8$, maintaining the number of parameters and FLOPs fixed. The SCHEME model with only BD-MLP improves on the baseline by 0.2%. The addition of the CCA branch provides an additional gain of 0.6%, showing the gains of better feature clusters. Since CCA is not used at inference, its gains are free in terms of additional parameters/FLOPs.
Effect of large expansion ratios: The right of Figure 1 shows the effect of simultaneously increasing the expansion ratio $E$ and adjusting groups to realize different models of similar size and complexity. All models are based on the Metaformer-PPAA-S12. For fixed parameters/FLOPs, SCHEMEformer models achieve a gain of 1.0% to 1.4% over the baseline by increasing $E$ from 1 to 4. The gains increase with larger expansions and saturate at larger FLOPs. This shows that higher internal feature dimensions are crucial for obtaining better accuracy with smaller ViT models.
Alternative designs of Channel Mixer: We investigate whether alternative choices to the CCA branch could accomplish this goal more effectively. Table 7 compares the performance of models that replace CCA with other feature mixing operations: the channel shuffling operation of ShuffleNet [64], a squeeze and excitation [23] network (SENet), a single layer of convolution, and a dynamic version of CCA (DyCCA), where the weight $\alpha$ of (6) is predicted dynamically, using GCT attention [35]. The table shows that CCA obtains the best result.
It is worth noting that while CCA has a lot more computation than some of these alternatives, it is not needed during inference. We found this not to be true for the alternatives, which produce much more balanced weights $\alpha$ after training convergence, and cannot be discarded at inference without a performance drop. We conjecture that, because all these alternatives have learnable parameters, the network learns to leverage them to extract complementary features, which must be fused with those produced by the BD-MLP branch during inference. Thus, while the alternative designs are very small modules, they incur an additional computational cost. CCA does not, achieving the best performance with no additional cost during inference.
[…]
@@
Regularizing effect of CCA: Fig. 3(a) shows the evolution of the weight $1-\alpha$ of the CCA branch in (6) during training. While initially large, it gradually decays to zero as training progresses. The figure shows that this holds for all network layers. Hence, CCA can be discarded at inference. Fig. 5 in appendix plots the weights $1-\alpha$ upon training convergence, for the family of SCHEMEformer-PPAA-44-e8 models, confirming that the weights are indeed very close to zero across all layers. Table 8 shows the impact of removing CCA at inference in top-1 accuracy. The performance decrease is negligible ($\approx 0.02$) across backbones.
[…]
(a) Evolution of weight $1-\alpha$
(b) Class separability
Figure 4: Regularizing effect of CCA. Left: Evolution of weight $1-\alpha$ across all layers for SCHEMEformer-PPAA-44-e8-S12 model. Right: Comparison of class separability of the features $\bf y$ of the SCHEME module, for 50 random classes on ImageNet-1K validation set, using SCHEMEformer-PPAA-44-e8-S12 model trained with and without CCA.
@@
We conjectured above that CCA helps training because it facilitates feature clustering into naturally independent groups that do not require inter-group communication. We tested this hypothesis by studying the intermediate feature vectors ${\bf v}_{l}$ (obtained after max-pooling the features $\bf y$ of (4)) extracted from four randomly selected layers $l$ of two ImageNet pretrained models, trained with and without CCA respectively. We split the features into the 4 groups used in the model ${\bf v}_{l,g},g\in\{1,\ldots 4\}$ and concatenated the features of all layers in the same group. This produced four vectors ${\bf u}_{g}=\mbox{concat}(\{v_{l,g}\}_{l})$ containing the features of each group $g$ extracted throughout the network. A linear classifier was then learned over each vector ${\bf u}_{g}$. Table 9 shows the top-1 accuracy per feature group and model. To evaluate whether groups learn different class clusters, we also average the outputs from the four group classifiers to obtain the final accuracy. Without CCA, i.e. no group communication during training, the network produces feature groups that tend to be individually more predictive of the image class, but less predictive when combined. This suggests that there is some amount of redundancy between the features extracted by the different groups. By introducing inter-group communication, CCA enables the groups to learn more diverse sets of features, that complement each other.
Fig. 3(b) shows the class separability of the intermediate features $\bf y$ of (4) of a randomly chosen layer of the SCHEMEformer-PPAA-44-e8-S12 model. Class separability was measured as in [37], with a final value obtained by averaging the class separability across all classes. The model trained with CCA has higher class separability than that without it. This confirms that CCA is helpful in forming feature clusters that increase class separability during training. Conversely, we tested if CCA is helpful when channel shuffling is inserted in between the two mixer MLP layers, which destroys the group structure. This variant of the SCHEMEformer-PPAA-44-e8-S12 model achieved an accuracy of 79.1% for both training with and without CCA. This shows that CCA is not helpful when feature groups are mixed. Similarly, CCA did not provide any gains when applied to the standard MLP branch with full feature mixing (one group). All these results suggest that CCA indeed helps to form the independent feature clusters needed to achieve the computational efficiency of channel groups without performance degradation.
5 Conclusion
@@
In this work, we proposed the SCHEME module with sparse feature mixing for improving the performance of ViTs. This uses a weighted fusion of a BD-MLP branch, which abstracts existing MLPs with block diagonal structure, and a parameter-free CCA branch that helps to cluster features into groups during training. The CCA branch was shown to improve training but not be needed at inference. Experiments showed that it indeed improves the class separability of the internal feature representation of the BD-MLP branch, helping create feature clusters that are informative of the image class. The standard MLP in transformer block is replaced with the SCHEME module to obtain a new family of SCHEMEformer models that were shown to improve performance for classification, detection, and segmentation, for fixed parameters and FLOPs with favorable latency. SCHEME is shown to be effective for various ViT architectures and was also shown to provide a flexible way to scale models, always outperforming models with smaller MLP expansion ratios having the same complexity.
Acknowledgment
We acknowledge and thank the use of the Nautilus platform for some of the experiments discussed above.
Appendix
Appendix A Ablation Studies
CCA #Params FLOPs Top-1 Acc
Model Used  (M) (G) (%)
SCHEMEformer-PPAA-44-e8-S24 21.2 4.1 81.88
SCHEMEformer-PPAA-44-e8-S24 21.2 3.3 81.88
SCHEMEformer-PPAA-44-e8-S36 55.2 10 83.20
SCHEMEformer-PPAA-44-e8-S36 55.2 8.0 83.15
SCHEMEformer-PPAA-12-e8-S12 21.3 3.8 82.00
SCHEMEformer-PPAA-12-e8-S12 21.3 3.4 81.98
SCHEMEformer-PPAA-12-e8-S24 40.0 7.3 82.80
SCHEMEformer-PPAA-12-e8-S24 40.0 6.5 82.76
SCHEMEformer-PPAA-12-e8-S36 58.8 10.8 84.00
SCHEMEformer-PPAA-12-e8-S36 58.8 9.6 83.95
Table 10: Ablation studies on removing CCA branch during inference for the family of SCHEMEformer-PPAA models on ImageNet-1K dataset.
CCA Group1 Group2 Group3 Group4 Ensemble Model (%) (%) (%) (%) (%) SCHEMEformer-44-e8-S12 29.63 37.59 37.57 35.89 73.95 SCHEMEformer-44-e8-S12 27.89 37.64 30.02 37.67 75.04 Table 11: Ablation study on the formation of feature clusters in the BD-MLP branch of the SCHEME module. We train a linear classifier on top of the four feature groups extracted from the final MLP mixer of the transformer block of the network. The model trained with CCA forms feature clusters that learn diverse and complementary set of features that can obtain 1% higher validation accuracy than the model trained without CCA.
[…]
Backbone CCA RetinaNet 1× Mask R-CNN 1× Used #P $AP$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{b}_{S}$ $AP_{M}$ $AP_{L}$P $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$ SCHEMEformer-44-e8-S12 21 38.3 58.0 40.4 21.0 41.4 52.3 31 39.8 61.9 42.9 24.1 53.0 42.3 SCHEMEformer-44-e8-S12 21 38.3 58.0 40.4 21.0 41.4 52.3 31 39.8 61.9 42.9 24.1 53.0 42.3 SCHEMEformer-44-e8-S24 31 38.8 58.7 41.2 22.5 41.5 53.5 41 40.9 62.5 44.6 24.6 55.6 43.8 SCHEMEformer-44-e8-S24 31 38.8 58.7 41.2 22.5 41.5 53.5 41 40.9 62.5 44.6 24.6 55.6 43.8 Table 13: Ablation study of removing CCA for COCO-17 Object Detection and Instance Segmentation. Removing CCA at inference does not impact the AP values as they are identical to the model using CCA at inference. $AP^{b}$ and $AP^{m}$ denote bounding box AP and mask AP, respectively.
CCA Semantic FPN #Par FLOPs mIoU (%) 15.5 36.4 40.9 15.5 34.3 40.9 24.8 49.8 42.5 24.8 45.7 42.5 Table 14: Ablation study of removing CCA for Semantic Segmentation results on ADE20K dataset.
@@ Appendix A Ablation Studies
Peak learning rate used in the paper $4\times 10^{-4}$
@@
A.1 Plot of the learned $1-\alpha$ weights in the SCHEME module for SCHEMEformer models
@@
Figure 5: Plot of the learned weight (1-$\alpha$) values across different layers of a network for the family of SCHEMEformer models. The weights reach a peak value near the middle of the network. We demonstrate that these peak weights are not yet converged and training the network for more epochs decays these weights to zero while also improving the accuracy. For example, training the SCHEMEformer-PPAA-12-e8-S12 model for 200 additional epochs reduced the weight norm of the vector of $1-\alpha$ weights from 0.18 to 0.05 showing that these weights gradually approach zero as the training progresses while improving the accuracy further by 0.3%.
@@
The learned weights $1-\alpha$ for the family of SCHEMEformer models is shown in Fig. 5. Interestingly, the learned weights coarsely approximate the shape of a gaussian distribution. The learned weights reach a peak value in the middle layers of the network and drop to zero for all the other layers. The middle of the network typically correspond to the initial few layers of the third stage of the model that contains the maximum number of transformer blocks for all the models shown in Fig. 5. We conjecture that the weights for these layers have not fully converged and that feature mixing can still be useful for these layers and so training for more epochs will allow the $1-\alpha$ weights of these layers to converge to zero. To test this hypothesis, we trained the SCHEMEformer-PPAA-12-e8-S12 model for an additional 200 epochs beyond the standard 300 epochs and observed that the peak value of the $1-\alpha$ weights decreased further by 0.11 as compared to the model trained for 300 epochs and the accuracy improved by 0.3%. The weight norm (of all layers) decreased from 0.18 for 300 epoch model to 0.05 for 500 epoch model. This confirms our hypothesis and training SCHEMEformer models for larger epochs can further improve the accuracy.
A.2 More results of removing CCA at inference
Table 10 shows the results of the full family of SCHEMEformer models with and without using CCA at inference. It is straightforward to observe that removing CCA does not impact the accuracy consistently across different model depths and larger models have slightly higher accuracy drop as compared to smaller models while still being negligible overall ($\leq 0.05\%$).

### v4

@@ 4.2 Ablation Studies
Contribution of BD-MLP and CCA branch: Table 5 shows an ablation of the contribution of the BD-MLP and CCA branches. Starting from the Metaformer-PPAA-11-e2-S12, with expansion ratio $E=2$ and dense MLP, we replace the channel mixer by SCHEME to obtain the SCHEMEformer-PPAA-44-e8-S12 ($g_{1}$=$g_{2}$=4 and $E$=8), which maintains the number of parameters and FLOPs constant. The SCHEME model with only BD-MLP improves on the baseline by 0.2%. The addition of the CCA branch provides an additional gain of 0.6%, showing the gains of better feature clusters. Since CCA is not used at inference, its gains are free in terms of additional parameters/FLOPs.
@@
Regularizing effect of CCA: Fig. 4 shows the evolution of the weight $1-\alpha$ of the CCA branch in (6) during training. While initially large, it gradually decays to zero as training progresses. This holds for all network layers. Hence, CCA can be discarded at inference. Fig. 1 in supplementary plots the weights $1-\alpha$ upon training convergence, for the family of SCHEMEformer-PPAA-44-e8 models, confirming that the weights are indeed very close to zero across all layers.
Effect of large expansion ratios: The left of Figure 4 shows the effect of simultaneously increasing the expansion ratio $E$ and adjusting groups to realize different models of similar size and complexity. All models are based on the Metaformer-PPAA-S12. For fixed parameters/FLOPs, SCHEMEformer models achieve a gain of 1.0% to 1.4% over the baseline by increasing $E$ from 1 to 4. The gains increase with larger expansions and saturate at larger FLOPs. This shows that higher internal feature dimensions are important for obtaining better accuracy with smaller ViTs.
Alternative designs of Channel Mixer: We investigate whether alternative choices to the CCA branch could accomplish this goal more effectively. Table 5 compares models that replace CCA with other feature mixing operations: the channel shuffling operation of ShuffleNet [60], a squeeze and excitation [20] network (SENet), a single layer of convolution, and a dynamic version of CCA (DyCCA), where the weight $\alpha$ of (6) is predicted dynamically, using GCT attention [30]. CCA obtains the best result. While CCA is computationally heavier than some of these alternatives, it is not needed at inference, as shown in Table 3. This is not true for the alternatives, which produce much more balanced weights $\alpha$ after training convergence, and cannot be discarded at inference without performance drop. We conjecture that their learnable parameters extract complementary features which must be used at inference.
@@
Feature Clustering: We conjectured above that CCA helps training because it facilitates feature clustering into naturally independent groups that do not require inter-group communication. We tested this hypothesis by studying the intermediate feature vectors ${\bf v}_{l}$ (obtained after max-pooling the features $\bf y$ of (4)) extracted from four randomly selected layers $l$ of two ImageNet pretrained models, trained with and without CCA. We split the features into the 4 groups used in the model ${\bf v}_{l,g},g\in\{1,\ldots 4\}$ and concatenated the features of all layers in the same group. This produced four vectors ${\bf u}_{g}=\mbox{concat}(\{v_{l,g}\}_{l})$ containing the features of each group $g$ extracted throughout the network. A linear classifier was then learned over each vector ${\bf u}_{g}$. Table 5 shows the top-1 accuracy per feature group and model. To evaluate whether groups learn different class clusters, we also average the outputs from the four group classifiers to obtain the final accuracy. Without CCA, i.e. no group communication during training, the network produces feature groups individually more predictive of the image class, but less predictive when combined. This suggests that there is redundancy between the features of the different groups. By introducing inter-group communication, CCA enables the groups to learn more diverse features, that complement each other. Fig. 4 shows the class separability of the intermediate features $\bf y$ of (4) of a randomly chosen layer of the SCHEMEformer-PPAA-44-e8-S12. Class separability was measured as in [32], with a final value obtained by averaging the class separability across all classes. The model trained with CCA has higher class separability than that without it. This confirms that CCA is helpful in forming feature clusters that increase class separability during training. Conversely, we tested if CCA is helpful when channel shuffling is inserted in between the two mixer MLP layers, which destroys the group structure. This variant of the SCHEMEformer-PPAA-44-e8-S12 model achieved an accuracy of 79.1% for both training with and without CCA (Table 5, row 1). This shows that CCA is not helpful when feature groups are mixed. Similarly, CCA did not provide any gains when applied to the standard MLP branch with full feature mixing. These results suggest that CCA indeed helps to form the independent feature clusters needed to achieve the computational efficiency of channel groups without performance degradation. Table 5 shows that, despite the higher accuracy of the individual group features of the model trained without CCA, the model trained with CCA has 0.6% higher ensemble accuracy.
5 Limitation and Conclusion
Although SCHEME module improves the accuracy-throughput curves of popular ViTs, it incurs a slight overhead in memory (see section A.3.2) during training due to the channel attention operation which limits the batch size on smaller GPUs. However, Figure 4 shows that the contribution of the CCA branch becomes negligible after 150 epochs. Therefore, the CCA branch can be removed beyond this point, significantly reducing the training overhead.
@@ 5 Limitation and Conclusion
In this work, we proposed the SCHEME module for improving the performance of ViTs. SCHEME leverages a block diagonal feature mixing structure to enable MLPs with larger expansion ratios, a property that is shown to improve transformer performance, without increase of model parameters or computation. During training, it uses a weighted fusion of a BD-MLP branch, and a parameter-free CCA branch that helps to cluster features into groups. The CCA branch was shown to improve training but not be needed at inference. Experiments showed that it indeed improves the class separability of the internal feature representation of the BD-MLP branch, helping create feature clusters that are informative of the image class. The standard transformer MLP was replaced with the SCHEME module to obtain a new family of SCHEMEformer models with improved performance for classification, detection, and segmentation, for fixed parameters and FLOPs with favorable latency. SCHEME was shown to be effective for various ViT architectures, always outperforming similar complexity models with smaller MLP expansion ratios.
Appendix A Appendix
Table 9: Comparison with state-of-the-art ViT models on the ImageNet-1K dataset. The SCHEME module enhances the accuracy of existing ViTs while maintaining or achieving higher throughput. * denotes the results reported in the original paper at 224 input resolution.
Model #Par (M) FLOPs (G) T (img/s) Acc (%)
ViT-Base [11] 86 17.6 112 79.7
SCHEME-ViT-12-e8-Base 77.1 15.5 130 (+16%) 79.9 (+0.2%)
DeiT-Tiny [37] 6 1.3 117 74.5
SCHEME-DeiT-12-e8-Tiny 7.5 1.6 117 (+0%) 76.0 (+1.5%)
Poolformer-S12 12.0 1.8 166 77.2
SCHEME-Poolformer-12-e8-S12 16.7 2.6 171 (+3%) 78.5 (+1.3%)
Metaformer-S12 12.0 1.8 133 78.9
SCHEMEformer-PPAA-12-e8-S12 11.8 1.8 133 (+0%) 79.7 (+0.8%)
CAformer-S12 25 4.2 74 82.9
SCHEME-CAformer-12-e8-S12 23.9 3.6 80 (+8%) 82.9 (+0.0%)
CoAtNet-0 25.0 4.2 88 81.6
SCHEME-CoAtNet-12-e8-0 24.0 4.1 110 (+25%) 81.6 (+0.0%)
CSWin-Tiny 23.0 4.3 20 82.8
SCHEME-CSWin-12-e8-Tiny 29.1 5.6 21 (+5%) 83.2 (+0.4%)
DaViT-Tiny 23.0 4.3 61 82.8
SCHEME-DaViT-12-e8-Tiny 37.0 6.6 62 (+1%) 83.0 (+0.2%)
T2T-ViT-14 21.5 6.1 70 81.7
SCHEME-T2T-ViT-12-e8-14 27.7 8.1 71 (+1%) 82.1 (+0.4%)
BiFormer-Tiny[61] 13 2.2 57 81.4
SCHEME-BiFormer-12-e3-Tiny 11.5 1.9 59 (+4%) 81.4 (+0.0%)
ViT-Large[11] 304 61.6 56 76.5*
SCHEME-ViT-44-e8-Large 143 28.7 114 (+104%) 81.7 (+5.2%)
Pyramid-Vision-Llama-S[3] 26 3.3 7 81.6
SCHEME-Pyramid-Vision-Llama-S 27 3.8 7 (+ 0%) 82.4 (+0.8%)
A.1 Code Release
Code is included in the supplementary material. Code and trained models will be released publicly upon acceptance of the paper.
A.2 Implementation Details
A.2.1 Hyperparameter Settings
Table A.2.1 shows the detailed hyperparameter settings of the family of SCHEMEformer models reported in the main paper. For reference, SCHEMEformer-44-e8-224 model requires about 3 days to train on 8 NVIDIA TITAN Xp GPUs.
@@ A.2.1 Hyperparameter Settings
Figure 6: Plot of the learned weight (1-$\alpha$) values across different layers of a network for the family of SCHEMEformer models. The weights reach a peak value near the middle of the network. We demonstrate that these peak weights are not yet converged and training the network for more epochs decays these weights to zero while also improving the accuracy. For example, training the SCHEMEformer-PPAA-12-e8-S12 model for 200 additional epochs reduced the weight norm of the vector of $1-\alpha$ weights from 0.18 to 0.05 showing that these weights gradually approach zero as the training progresses while improving the accuracy further by 0.4%.
A.3 Ablation Studies
A.3.1 Plot of the learned 1-$\alpha$ weights in the SCHEME module for SCHEMEformer models
@@
The learned weights 1-$\alpha$ for the SCHEMEformer model family are shown in Fig. 6. Interestingly, the learned weights coarsely approximate the shape of a gaussian distribution. The learned weights reach a peak value in the middle layers of the network and drop to zero for all the other layers. The middle of the network typically corresponds to the initial few layers of the third stage of the model that contains the maximum number of transformer blocks for all the models shown in Fig. 6. We conjecture that the weights for these layers have not fully converged and that feature mixing can still be useful for these layers and so training for more epochs will allow the 1-$\alpha$ weights of these layers to converge to zero. To test this hypothesis, we trained the SCHEMEformer-PPAA-12-e8-S12 model for an additional 200 epochs beyond the standard 300 epochs and observed that the peak value of the 1-$\alpha$ weights decreased further by 0.11 as compared to the model trained for 300 epochs and the accuracy improved by 0.4%. The weight norm (of all layers) decreased from 0.18 for 300 epoch model to 0.05 for 500 epoch model. This confirms our hypothesis and training SCHEMEformer models for larger epochs can further improve the accuracy.
Table 11: Training Overhead of CCA. CCA adds only a small overhead in GPU memory and training time.
CCA #Par Train FLOPs Val Acc GPU Mem. Train Throughput
Model (M)$\downarrow$ (G)$\downarrow$ (%) $\uparrow$ (G) $\downarrow$ (iters/s) $\uparrow$
SCHEMEformer-PPAA-44-e8-S12 11.8 1.77 79.1 8 215
SCHEMEformer-PPAA-44-e8-S12 11.8 2.16 79.7 9 180
Table 12: Ablation study of longer training for SCHEMEformer-PPAA-44-e8-S12.
Model #P (M) FLOPs (G) Throughput (img/s) 300-Acc (%) 500-Acc (%)
Metaformer-11-e2-S12 (Baseline) 11.8 1.77 133 78.9 79.6
SCHEMEformer-44-e8-S12 11.8 1.77 133 79.7 80.1
A.3.2 Training Overhead of CCA
Table A.3.1 compares the training GPU memory and throughput for SCHEME mixer with and without using CCA. CCA improves the accuracy by 0.6% with only a slight increase in the GPU memory (+12.5%) and training time (+16.7%). Further, CCA is not needed during inference thereby providing gains for "free" without additional computational cost at inference.
A.3.3 Training for longer epochs
Table A.3.1 shows the comparison of training for longer epochs for SCHEMEformer with the baseline Metaformer model. We train for an additional 200 epochs from the standard 300 epochs. SCHEMEformer-PPAA-44-e8-S12 trained for 300 epochs even outperforms the baseline model trained for 500 epochs. On continuing the training from 300 to 500 epochs, SCHEMEformer continues to improve the performance without saturation suggesting that it is beneficial to train with CCA for longer epochs. 300 to 500 epochs is a much larger increase of training time (67%) than the 16.7% increase in training time required by SCHEME mixer (see Table A.3.1).
Table 13: Impact of removing CCA branch during inference.
CCA #Params FLOPs Top-1 Acc
Model Used  (M) (G) (%)
SCHEMEformer-PPAA-44-e8-S12 11.83 2.16 79.74
SCHEMEformer-PPAA-44-e8-S12 11.83 1.77 79.72
SCHEMEformer-PPAA-12-e8-S24 40.0 7.3 82.80
SCHEMEformer-PPAA-12-e8-S24 40.0 6.5 82.76
SCHEMEformer-PPAA-12-e8-S36 58.8 10.8 84.00
SCHEMEformer-PPAA-12-e8-S36 58.8 9.6 83.95
SCHEME-CoatNet-44-e8-0 17.80 3.83 80.70
SCHEME-CoatNet-44-e8-0 17.80 3.42 80.68
SCHEME-Swin-12-e8-T 36.93 7.00 81.69
SCHEME-Swin-12-e8-T 36.93 5.89 81.69
Table 14: Ablation study of removing CCA for COCO-17 Object Detection and Instance Segmentation. Removing CCA at inference does not impact the AP values as they are identical to the model using CCA at inference. $AP^{b}$ and $AP^{m}$ denote bounding box AP and mask AP, respectively. Backbone models denote SCHEMEformer-PPAA-44-e8- variants.
Backbone CCA RetinaNet 1× Mask R-CNN 1×
Used #P $AP$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{b}_{S}$ $AP_{M}$ $AP_{L}$P $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
S12 21 38.3 58.0 40.4 21.0 41.4 52.3 31 39.8 61.9 42.9 24.1 53.0 42.3
S12 21 38.3 58.0 40.4 21.0 41.4 52.3 31 39.8 61.9 42.9 24.1 53.0 42.3
S24 31 38.8 58.7 41.2 22.5 41.5 53.5 41 40.9 62.5 44.6 24.6 55.6 43.8
S24 31 38.8 58.7 41.2 22.5 41.5 53.5 41 40.9 62.5 44.6 24.6 55.6 43.8
Table 15: Ablation study of removing CCA for Semantic Segmentation results on ADE20K validation dataset.
CCA Semantic FPN
#Par FLOPs mIoU (%)
15.5 36.4 40.9
15.5 34.3 40.9
24.8 49.8 42.5
24.8 45.7 42.5
A.3.4 Impact of removing CCA at inference
Table A.3.3 shows the impact of removing CCA at inference, for various backbones. While the number of FLOPS decreases, the top-1 accuracy changes very little ($\approx 0.02$ difference). Hence, there is no advantage in using CCA at inference. This is unlike training, where the use of CCA makes a non-negligible difference, as shown in Table 6 of the main paper.